# thanks: $0$-0003-0728-2533thanks: $0$-0001-5501-6008

### v1

@@ III.1.1 Convergence tests for extremely cuspy subhalos
As our result for $\gamma=1.5$ differs significantly from that shown found by Ref. [51], we conducted additional convergence tests for this case. There are three parameters that control the time and special resolution of the simulation: (1) time step size criterion; (2) softening length, $r_{\rm softening}$; (3) particle number, $N$.
@@
As we decrease the parameter $\eta$, the results slowly converge, but the differences from the results of Ref. [51] remain. Possible explanations for the differences include: a) To run the simulations, [51] uses SUPERBOX [85] which computes the gravitational interactions using multiple layers of grids while we use Gadget-4, which computes the gravitational interactions using FFM; b) the details of our self-binding analysis differ from the approach in Ref. [51]  [^4] Even for the cases where the tidal tracks we find agree well with the results of [51], the dependence of $V_{\rm max}/V_{\rm max,0}$ and $R_{\rm max}/R_{\rm max,0}$ on $M_{\rm bound}/M_{\rm bound,0}$ differ from those found by [51], indicating some difference in which particles are considered to be bound to the subhalo. For example, for the case $(\alpha,\beta,\gamma)=(1,3,1)$, we find that we must multiply the bound mass we obtain by a factor of $2$, to obtain an approximate match with $V_{\mathrm{max}}$ as a function of $M_{\mathrm{bound}}/M_{\mathrm{bound},0}$ as found by [51], see Appendix A.
Interestingly, we find that using a global time step size, i.e. all particles are evolved using the same time step, significantly improves convergence allowing the use of a larger time step, see the right panel of Fig. 4. The computing time required to achieve converged results is then much shorter than that using individual time step sizes for each particle. This suggests that to simulate extremely cuspy subhalos, the time step size criterion should be chosen carefully and using a global time step size may be helpful (see also Ref. [86, 87]). We also find that, when we use a very large time step size, e.g. $\eta=0.5$, the results are unconverged, but are in fact closer to the fitting function obtained by [51]. For this case, the time step size $\Delta t=1.2\ {\mathrm{Myr}}$, which is still smaller than the one used by [51], i.e. $\Delta t=4.6\ {\mathrm{Myr}}$. Therefore, it is possible that the results for the case with $\gamma=1.5$ shown in [51] are not fully converged. A direct comparison between Gadget-4 and SUPERBOX using the same initial conditions and post-analysis would be required to further confirm this hypothesis.
@@ III.2 Evolution of density profiles
where the concentration parameter $c$ enters in the virial radius, and $r_{\mathrm{te}}$, $f_{\mathrm{t}}$ and $\delta$ are also concentration-dependent:
@@
Since, in the current work, we fix the subhalo concentration at $20.5$ (see Eqs. (11) and (12)), we choose to fit the density profile measured from our simulations using the simpler formula, Eq. (23). Ref. [52] found that their best fit value for $\delta$ is 2–3. Therefore, we consider two cases: $\delta=3$ and $\delta=2$. We emphasize that sometimes it can be useful to have a density profile that leads to analytic formula for certain physical quantities, such as the enclosed mass within a given radius. The case with $\delta=2$ has been widely used in modeling subhalo density profiles in the analysis of strong gravitational lenses (e.g. [34]). Applying the transfer function Eq. (23) with $\delta=2$ to the NFW profile, the lensing convergence can be computed analytically [89].
We first measure the density transfer function $H$ of subhalos at the apocenters of their orbits from our simulations with different initial subhalo profiles. Then we fit Eq. (23) to these measured transfer functions to find the best fit $r_{\mathrm{te}}$ and $r_{t}$. $r_{\mathrm{te}}$ and $r_{t}$ as a function of the bound mass fraction for different $\gamma$. Results are shown in Fig. 5 (colored markers). For each value of $\gamma$, we have combined the simulation data for different subhalo orbits, i.e. different value of $R_{\mathrm{p}}/R_{\mathrm{a}}$. Inspired by the work of Ref. [50] and Ref. [52], we fit the mass dependence in $r_{\mathrm{te}}$ and $r_{t}$ assuming
@@
The best fit formula are shown in Fig. 5 (colored curves) for $(\alpha,\beta)=(1,3)$ and different $\gamma$. We also compare our fitting functions to those obtained by Hayashi et al. [50] (black curve) and Green et al. [52] (grey curve) for initial NFW profiles, i.e. $\gamma=1$. We find that the effective tidal radius $r_{\mathrm{te}}$ we obtain is in broad agreement with that from Hayashi et al. at $x>4\times 10^{-2}$. At $x<4\times 10^{-2}$, the fitting formula from Hayashi et al. shows an upturn, which does not appear in our results. However, we note that the Hayashi et al. fitting function was calibrated only for $x>4\times 10^{-2}$. Our result for $r_{\mathrm{te}}$ is lower than that found by Green et al. Here we have taken into account the difference in definition of halo concentration in our work from Green et al., but this has only a small effect. Since the number of particles in our simulations is $10$ times higher than that used by Green et al., we can better resolve the central region of the subhalo. We have tried excluding radii bins with $r<10^{-2}r_{\mathrm{vir,sub}}$ from our fitting. Our results are then in good agreement with those from Green et al., as shown by the stars in Fig. 5. This suggests that the difference we see is due to numerical resolution and the range of radii bins used in the fitting process. For $f_{\mathrm{t}}$, a similar difference between our results and those obtained by Green et al. is also seen, as shown in the right panel of Fig. 5. The fitting function from Green et al. overestimates the density decrease in the central region of subhalos due to tidal effects. Similar findings has been reported in Ref. [53]. Again, we show that we can recover the Green et al. results by excluding the data points at small radii from the our fitting procedure.
Figure 5: Best-fit parameters for the density profile as a function of bound mass fraction for $(\alpha,\beta)=(1,3)$ and different $\gamma$. Left: the effective tidal radius in Eq. (23). The filled (open) circles are obtained by fitting Eq. (23) to the density profiles measured from simulation assuming $\delta=3$ ($\delta=2$). For each $\gamma$, data from simulations with different $R_{\mathrm{p}}/R_{\mathrm{a}}$ are combined. The colored lines show the fitting formula Eq. (30) with our best-fit parameter values. The fitting functions from Hayashi et al. 2003 [50] (black) and Green et al. 2019 [52] (grey) are also shown. Right: the normalization parameter $f_{\mathrm{t}}$ in Eq (23). The purple stars show the best-fit values for $r_{\mathrm{te}}$ and $f_{\mathrm{t}}$ when excluding measurements of the density profiles at $r<10^{-2}$. The Hayashi et al. 2003 fitting function is calibrated only for $M_{\mathrm{bound}}/M_{\mathrm{bound,0}}>4\times 10^{-2}$, thus at small mass ratios an extrapolation has be used (black dotted line).
@@
Table 2: Best-fit parameters in Eqs. (30) and (31) for density transfer function Eq. 23 with different initial profiles. $\delta=3$ has been assumed.
@@
Table 3: Best-fit parameters in Eqs. (30) and (31) for density transfer function Eq. 23 with different initial profiles. $\delta=2$ has been assumed.
@@ IV Semi-analytic models
In the remainder of this section we detail the orbital and tidal physics included in our model, and present results for the calibration of the parameters of this model. All the models described below are implemented in the public semi-analytic model for galaxy formation, Galacticus [82] [^5] https://github.com/galacticusorg/galacticus.
@@ IV.1 Orbital evolution
where $\mathbf{a}_{\mathrm{g}}$ is the gravitational acceleration from the host, and $\mathbf{a}_{\mathrm{df}}$ is the dynamical friction caused by the overdense wake of host particles that generated behind the subhalo when it orbits within the host. Using the Chandrasekhar formula, $\mathbf{a}_{\mathrm{df}}$ can be computed as [94]:
@@ IV.2 Tidal stripping
where $\alpha_{\mathrm{s}}$ is a free parameter that controls the efficiency of tidal stripping. There exist several relevant physical timescales that could be chosen for $T_{\mathrm{loss}}$: a) orbital time scale: $T_{\rm orbit}=2\pi/\max\{\omega_{\mathrm{t}},\omega_{\mathrm{r}}\}$ with $\omega_{\mathrm{t}}$ and $\omega_{\mathrm{r}}$ are the instantaneous frequencies of tangential and radial motion, respectively (this choice is adopted by Yang et al. 91); b) the dynamical time scale $T_{\rm dyn}=\sqrt{3\pi r_{\mathrm{t}}^{3}/16\mathrm{G}M_{\rm sub}(<r_{\mathrm{t}})}$. These two timescales are related but may differ significantly when the subhalo is close to the pericenter of its orbit. In this work, we find that the latter results in better fits to measurements from our simulations. Thus we will use it as our fiducial choice.
@@ IV.3 Tidal heating
As a subhalo orbits within its host and loses mass due to tidal stripping, its density profile also changes with time due to two effects: a) after some mass is removed from the subhalo, it will revirialize and approach a new equilibrium with a different density profile; b) particles in the subhalo gain energy from tidal shocks, a.k.a. the tidal heating effect, leading to the expansion of subhalo. When the subhalo passes through the pericenter of its orbit, both tidal stripping and tidal heating are strong. The subhalo loses a large fraction of the mass outside the tidal radius and at the same time is heated. As the subhalo approaches the apocenter of its orbit, it will begin to revirialize, resulting in a less concentrated density profile. As the subhalo becomes less concentrated, the tidal radius shrinks further and more mass will be stripped from the subhalo. The process of revirialization is complicated [104]. In the current work, we focus on the density profile at successive apocenters for two reasons. First, the subhalo spends more time near apocenter during its orbit through its host. Second, at apocenter, the subhalos have had enough time since the strong mass loss and heating at pericenter to be revirialized, allowing us to apply the heating model proposed in [59, 61].
@@
where $M_{\mathrm{i}}$ is the enclosed mass within $r_{\mathrm{i}}$. Note that if no shell crossing happens, $M_{\mathrm{i}}$ is unchanged after expansion by definition. Given the initial density of the mass shell, $\rho_{\mathrm{i}}$, and the relation between $r_{\mathrm{i}}$ and $r_{\mathrm{f}}$ from solving equation (36), the density of the mass shell after reaching new equilibrium can be written as
@@
where $\epsilon_{\mathrm{h}}$ is a coefficient needed to be calibrated to simulations, $\omega_{\mathrm{p}}$ is the angular frequency of particles at the half-mass radius of the subhalo, $T_{\rm shock}=r_{\rm sub}/V_{\rm sub}$ is time scale of tidal shock, $\gamma_{\mathrm{h}}$ is the adiabatic index, $g_{ab}$ is the tidal tensor, and $G_{ab}$ is the time integral of $g_{ab}$ [61, 91]:
@@
The term in the square brackets in Eq. (38) accounts for the adiabatic correction, i.e. the heating energy gained by particles with orbital timescale much smaller than $T_{\rm shock}$ is suppressed due to ‘adiabatic shielding’ [105, 106, 107]. Gnedin et al.  [104] find $\gamma_{\mathrm{h}}=2.5$ for $\omega_{\mathrm{p}}T_{\rm shock}\lesssim 1$, while it is shown that $\gamma_{\mathrm{h}}=1.5$ in the regime $\omega_{\mathrm{p}}T_{\rm shock}\gtrsim 4$. On the other hand, [91] find that a value of $\gamma_{\mathrm{h}}=0$ predicts a $V_{\mathrm{max}}-M_{\mathrm{bound}}$ relation that is in better agreement with high-resolution cosmological N-body simulations Caterpillar [93] and ELVIS [92]. In this work, we have tried both $\gamma_{\mathrm{h}}=2.5$ and $\gamma_{\mathrm{h}}=0$, and reach a similar conclusion as [91], i.e. $\gamma_{\mathrm{h}}=0$ results in $V_{\mathrm{max}}$ matching more closely with simulation results. This is partially due to the fact that the decaying term introduced in Eq. (39) also acts to suppress the heating energy.
@@
where $\epsilon_{1}$ and $\epsilon_{2}$ are the contributions from the first-order and second-order energy perturbations, $f_{2}$ is a coefficient, $\chi_{v}$ is the position-velocity correlation (see Ref. [104]), and $\sigma_{r}$ is the radial velocity dispersion of subhalo prior to any tidal heating. Ref. [104] show that $\chi_{v}$ depends only weakly on the density profile, so we fix $\chi_{v}$ at a typical value of $-0.333$. Any uncertainty of $\chi_{v}$ is absorbed in the coefficient $f_{2}$.
@@
To ensure that, after tidal heating, the density profile remains continuous at $r_{\rm crossing}$ [^6] From Eqs. (36) and (37), this requires that $\Delta\epsilon$ is differentiable., we require that both $\xi$ and $\mathrm{d}\xi/\mathrm{d}r$ are continuous at $r_{\rm crossing}$ or equivalently

### v2

@@ III.1.1 Convergence tests for extremely cuspy subhalos
As our result for $\gamma=1.5$ differs significantly from that shown found by Ref. [53], we conducted additional convergence tests for this case. There are three parameters that control the time and spatial resolution of the simulation: (1) time step size criterion; (2) softening length, $r_{\rm softening}$; (3) particle number, $N$.
@@
As we decrease the parameter $\eta$, the results slowly converge, but the differences from the results of Ref. [53] remain. Possible explanations for the differences include the following: (i) To run the simulations, [53] uses SUPERBOX [92] which computes the gravitational interactions using multiple layers of grids while we use Gadget-4, which computes the gravitational interactions using FFM; (ii) the details of our self-binding analysis differ from the approach in Ref. [53] [^4] Even for the cases where the tidal tracks we find agree well with the results of [53], the dependence of $V_{\rm max}/V_{\rm max,0}$ and $R_{\rm max}/R_{\rm max,0}$ on $M_{\rm bound}/M_{\rm bound,0}$ differ from those found by [53], indicating some difference in which particles are considered to be bound to the subhalo. For example, for the case $(\alpha,\beta,\gamma)=(1,3,1)$, we find that we must multiply the bound mass we obtain by a factor of $2$, to obtain an approximate match with $V_{\mathrm{max}}$ as a function of $M_{\mathrm{bound}}/M_{\mathrm{bound},0}$ as found by [53], see Appendix A
Interestingly, we find that using a global time step size, i.e. all particles are evolved using the same time step, significantly improves convergence allowing the use of a larger time step, see the right panel of Fig. 4. The computing time required to achieve converged results is then much shorter than that using individual time step sizes for each particle. This suggests that to simulate extremely cuspy subhalos, the time step size criterion should be chosen carefully and using a global time step size may be helpful (see also Refs. [93, 94]). We also find that, when we use a very large time step size, e.g. $\eta=0.5$, the results are unconverged, but are in fact closer to the fitting function obtained by [53]. For this case, the time step size $\Delta t=1.2\ {\mathrm{Myr}}$, which is still smaller than the one used by [53], i.e. $\Delta t=4.6\ {\mathrm{Myr}}$. Therefore, it is possible that the results for the case with $\gamma=1.5$ shown in [53] are not fully converged. A direct comparison between Gadget-4 and SUPERBOX using the same initial conditions and post-analysis would be required to further confirm this hypothesis.
@@ III.2 Evolution of density profiles
where the concentration parameter $c$ enters in the virial radius, and $r_{\mathrm{te}}$, $f_{\mathrm{t}}$ and $\delta$ are also concentration dependent:
@@
Since, in the current work, we fix the subhalo concentration at $20.5$ [see Eqs. (11) and (12)], we choose to fit the density profile measured from our simulations using the simpler formula, Eq. (23). Reference [54] found that their best fit value for $\delta$ is 2–3. Therefore, we consider two cases: $\delta=3$ and $\delta=2$. We emphasize that sometimes it can be useful to have a density profile that leads to analytic formula for certain physical quantities, such as the enclosed mass within a given radius. The case with $\delta=2$ has been widely used in modeling subhalo density profiles in the analysis of strong gravitational lenses (e.g. [35]). Applying the transfer function Eq. (23) with $\delta=2$ to the NFW profile, the lensing convergence can be computed analytically [96].
We first measure the density transfer function $H$ of subhalos at the apocenters of their orbits from our simulations with different initial subhalo profiles. Then we fit Eq. (23) to these measured transfer functions to find the best fit $r_{\mathrm{te}}$ and $r_{t}$. $r_{\mathrm{te}}$ and $r_{t}$ as a function of the bound mass fraction for different $\gamma$. Results are shown in Fig. 5 (colored markers). For each value of $\gamma$, we have combined the simulation data for different subhalo orbits, i.e. different value of $R_{\mathrm{p}}/R_{\mathrm{a}}$. Inspired by the work of Refs. [52, 54], we fit the mass dependence in $r_{\mathrm{te}}$ and $r_{t}$ assuming
@@
The best fit formula are shown in Fig. 5 (colored curves) for $(\alpha,\beta)=(1,3)$ and different $\gamma$. We also compare our fitting functions to those obtained by Hayashi et al. [52] (black curve) and Green et al. [54] (gray curve) for initial NFW profiles, i.e. $\gamma=1$. We find that the effective tidal radius $r_{\mathrm{te}}$ we obtain is in broad agreement with that from Hayashi et al. at $x>4\times 10^{-2}$. At $x<4\times 10^{-2}$, the fitting formula from Hayashi et al. shows an upturn, which does not appear in our results. However, we note that the Hayashi et al. fitting function was calibrated only for $x>4\times 10^{-2}$. Our result for $r_{\mathrm{te}}$ is lower than that found by Green et al. Here we have taken into account the difference in definition of halo concentration in our work from Green et al., but this has only a small effect. Since the number of particles in our simulations is $10$ times higher than that used by Green et al., we can better resolve the central region of the subhalo. We have tried excluding radii bins with $r<10^{-2}r_{\mathrm{vir,sub}}$ from our fitting. Our results are then in good agreement with those from Green et al., as shown by the stars in Fig. 5. This suggests that the difference we see is due to numerical resolution and the range of radii bins used in the fitting process. For $f_{\mathrm{t}}$, a similar difference between our results and those obtained by Green et al. is also seen, as shown in the right panel of Fig. 5. The fitting function from Green et al. overestimates the density decrease in the central region of subhalos due to tidal effects. Similar findings has been reported in Ref. [55]. Again, we show that we can recover the Green et al. results by excluding the data points at small radii from the our fitting procedure.
Figure 5: Best-fit parameters for the density profile as a function of bound mass fraction for $(\alpha,\beta)=(1,3)$ and different $\gamma$. Left: the effective tidal radius in Eq. (23). The filled (open) circles are obtained by fitting Eq. (23) to the density profiles measured from simulation assuming $\delta=3$ ($\delta=2$). For each $\gamma$, data from simulations with different $R_{\mathrm{p}}/R_{\mathrm{a}}$ are combined. The colored lines show the fitting formula Eq. (30) with our best-fit parameter values. The fitting functions from Hayashi et al. 2003 [52] (black) and Green et al. 2019 [54] (gray) are also shown. Right: the normalization parameter $f_{\mathrm{t}}$ in Eq (23). The purple stars show the best-fit values for $r_{\mathrm{te}}$ and $f_{\mathrm{t}}$ when excluding measurements of the density profiles at $r<10^{-2}$. The Hayashi et al. 2003 fitting function is calibrated only for $M_{\mathrm{bound}}/M_{\mathrm{bound,0}}>4\times 10^{-2}$, thus at small mass ratios an extrapolation has be used (black dotted line).
@@
Table 2: Best-fit parameters in Eqs. (30) and (31) for density transfer function Eq. (23) with different initial profiles. $\delta=3$ has been assumed.
@@
Table 3: Best-fit parameters in Eqs. (30) and (31) for density transfer function Eq. (23) with different initial profiles. $\delta=2$ has been assumed.
@@ IV Semi-analytic models
In the remainder of this section we detail the orbital and tidal physics included in our model, and present results for the calibration of the parameters of this model. All the models described below are implemented in the public semi-analytic model for galaxy formation, Galacticus [86] [^5] https://github.com/galacticusorg/galacticus
@@ IV.1 Orbital evolution
where $\mathbf{a}_{\mathrm{g}}$ is the gravitational acceleration from the host, and $\mathbf{a}_{\mathrm{df}}$ is the dynamical friction caused by the overdense wake of host particles that generated behind the subhalo when it orbits within the host. Using the Chandrasekhar formula, $\mathbf{a}_{\mathrm{df}}$ can be computed as [99]
@@ IV.2 Tidal stripping
where $\alpha_{\mathrm{s}}$ is a free parameter that controls the efficiency of tidal stripping. There exist several relevant physical time scales that could be chosen for $T_{\mathrm{loss}}$: (i) orbital time scale: $T_{\rm orbit}=2\pi/\max\{\omega_{\mathrm{t}},\omega_{\mathrm{r}}\}$ with $\omega_{\mathrm{t}}$ and $\omega_{\mathrm{r}}$ are the instantaneous frequencies of tangential and radial motion, respectively (this choice is adopted by Yang et al. 98); (ii) the dynamical time scale $T_{\rm dyn}=2\pi\sqrt{r_{\mathrm{t}}^{3}/16\mathrm{G}M_{\rm sub}(<r_{\mathrm{t}})}$ [108]. These two time scales are related but may differ significantly when the subhalo is close to the pericenter of its orbit. In this work, we find that the latter results in better fits to measurements from our simulations. Thus we will use it as our fiducial choice.
@@ IV.3 Tidal heating
As a subhalo orbits within its host and loses mass due to tidal stripping, its density profile also changes with time due to two effects: (i) after some mass is removed from the subhalo, it will revirialize and approach a new equilibrium with a different density profile; (ii) particles in the subhalo gain energy from tidal shocks, also known as the tidal heating effect, leading to the expansion of subhalo. When the subhalo passes through the pericenter of its orbit, both tidal stripping and tidal heating are strong. The subhalo loses a large fraction of the mass outside the tidal radius and at the same time is heated. As the subhalo approaches the apocenter of its orbit, it will begin to revirialize, resulting in a less concentrated density profile. As the subhalo becomes less concentrated, the tidal radius shrinks further and more mass will be stripped from the subhalo. The process of revirialization is complicated [109]. In the current work, we focus on the density profile at successive apocenters for two reasons. First, the subhalo spends more time near apocenter during its orbit through its host. Second, at apocenter, the subhalos have had enough time since the strong mass loss and heating at pericenter to be revirialized, allowing us to apply the heating model proposed in [61, 63].
@@
where $M_{\mathrm{i}}$ is the enclosed mass within $r_{\mathrm{i}}$. Note that if no shell crossing happens, $M_{\mathrm{i}}$ is unchanged after expansion by definition. Given the initial density of the mass shell, $\rho_{\mathrm{i}}$, and the relation between $r_{\mathrm{i}}$ and $r_{\mathrm{f}}$ from solving Eq. (36), the density of the mass shell after reaching new equilibrium can be written as
@@
where $\epsilon_{\mathrm{h}}$ is a coefficient needed to be calibrated to simulations, $\omega_{\mathrm{p}}$ is the angular frequency of particles at the half-mass radius of the subhalo, $T_{\rm shock}=r_{\rm sub}/V_{\rm sub}$ is the time scale of tidal shock, $\gamma_{\mathrm{h}}$ is the adiabatic index, $g_{ab}$ is the tidal tensor, and $G_{ab}$ is the time integral of $g_{ab}$ [63, 98]:
@@
The term in the square brackets in Eq. (38) accounts for the adiabatic correction, i.e. the heating energy gained by particles with orbital timescale much smaller than $T_{\rm shock}$ is suppressed due to “adiabatic shielding” [110, 111, 112]. Gnedin et al.  [109] find $\gamma_{\mathrm{h}}=2.5$ for $\omega_{\mathrm{p}}T_{\rm shock}\lesssim 1$, while it is shown that $\gamma_{\mathrm{h}}=1.5$ in the regime $\omega_{\mathrm{p}}T_{\rm shock}\gtrsim 4$. On the other hand, [98] find that a value of $\gamma_{\mathrm{h}}=0$ predicts a $V_{\mathrm{max}}-M_{\mathrm{bound}}$ relation that is in better agreement with high-resolution cosmological $N$-body simulations Caterpillar [45] and ELVIS [44]. In this work, we have tried both $\gamma_{\mathrm{h}}=2.5$ and $\gamma_{\mathrm{h}}=0$, and reach a similar conclusion as [98], i.e. $\gamma_{\mathrm{h}}=0$ results in $V_{\mathrm{max}}$ matching more closely with simulation results. This is partially due to the fact that the decaying term introduced in Eq. (39) also acts to suppress the heating energy.
@@
where $\epsilon_{1}$ and $\epsilon_{2}$ are the contributions from the first-order and second-order energy perturbations, $f_{2}$ is a coefficient, $\chi_{v}$ is the position-velocity correlation (see Ref. [109]), and $\sigma_{r}$ is the radial velocity dispersion of subhalo prior to any tidal heating. Reference [109] shows that $\chi_{v}$ depends only weakly on the density profile, so we fix $\chi_{v}$ at a typical value of $-0.333$. Any uncertainty of $\chi_{v}$ is absorbed in the coefficient $f_{2}$.
@@
To ensure that, after tidal heating, the density profile remains continuous at $r_{\rm crossing}$ [^6] From Eqs. (36) and (37), this requires that $\Delta\epsilon$ is differentiable. we require that both $\xi$ and $\mathrm{d}\xi/\mathrm{d}r$ are continuous at $r_{\rm crossing}$ or equivalently